Question 2

Without using finance calculator,

(a) Haziq invests RM15,000 in a fund at 7% compounded quarterly. After t years, the value of the investment is RM300,000. Determine the value of t.

(b) Sarah invested RM5000 into an account that pays 5% compounded semiannually. She intended to keep the account untouched for 5 years. However, after 3 years, she had to withdraw RM3000. Find the amount left in the account after five years from the time she made her investment.
1
Expert's answer
2020-05-25T09:42:43-0400

a)

"300,000=15,000*1.07^n"

"20=1.07^n"

"1.07^{45}=21"

"1.07^{44}=19.63"

After 45 years

b)

"5,000*1.05^3=5,788.125"

"(5,788.125-3,000)*1.05^2=3,073.91"